# Asymptotic computations of tropical refined invariants in genus 0 and 1

### v1

@@ 1.1. Context
Given some points in the plane, the problem of determining how many curves of fixed degree and genus pass through these points is a well-known question, which generalizes to other surfaces. Given a non-singular surface $X$ equipped with a sufficiently ample line bundle $\mathcal{L}$, curves on $X$ may be obtained as zero-sets of sections of $\mathcal{L}$. For a non-negative integer $g$, how many such curves of genus $g$ pass through $c_{1}(X)\cdot\mathcal{L}-1+g$ points on $X$ We denote by $N_{g}^{X}(\mathcal{L})$ this number, which is also known as the degree of the corresponding Severi variety. As a consequence of the adjunction formula, we could consider a dual question: given $\delta$ a number of nodes, what is the number $N_{X}^{\delta}(\mathcal{L})$ of $\delta$-nodal curves passing through the appropriate number of points on $X$
@@ 1.1.3. Interpretations and applications of tropical refined invariants
    Bousseau proved [Bou19] that through the change of variable $q=e^{i\hbar}$, tropical refined invariants actually compute generating series of (log-)Gromov Witten invariants with a $\lambda$-class.
@@ 1.2. Overview of results
For $X$ a non-singular toric surface associated to a $h$-transverse and horizontal polygon, one has:
@@
We then give a result concerning generating series when summing over $i$: let $AR_{g}^{X}(\beta)=\sum_{i=0}^{\infty}AR^{X}_{g,i}(\beta)x^{i}$ be this series, which we call the genus $g$ asymptotic refined invariant. In [Mév23], the second author proved that $AR^{X}_{0}(\beta)=p(x)^{\chi}$. However, the method used there does not seem to be manageable to deal with higher genus. With a slightly different point of view, we give in Theorem 4.17 a different proof of this fact. The interest is that this point of view also allows to perform the computation in genus 1. A priori, the polynomiality behavior for general $h$-transverse toric surface has not been proven in [BJP22], but in the genus 0 and 1 case it actually follows from our computations.
@@
The form taken by generating series for fixed genus suggests a nice but subtle regularity of the asymptotic of the refined invariants. However, we are for now limited by the computational techniques at our disposal, as the complexity of the computations increases quite fast with the genus or the codegree. With some time, the techniques presented should be able to provide a result for genus $2$ as well as codegree $2$, but at the cost of many lengthy computations. Given the above results, we conjecture the following.
@@
The asymptotic refined invariant $AR^{X}_{g}$ has the following form:
@@
Conjecture 1.1 treats the regularity of the asymptotic invariant for a fixed surface $X$. Similarly to the Göttsche conjecture, it would be interesting to study the dependence of $AR_{g}^{X}$ in the surface $X$. Hopefully, the polynomials $AR_{g,i}^{X}(\beta)$ are actually given by a universal polynomial $Q_{g,i}(\beta^{2},\beta\cdot K_{X},K_{X}^{2},\chi(X))$. Theorems 3.2 and 5.18 prove that it is the case for $g=0,1$ and $i=0,1$.
@@ 1.3. Organization of the paper
Section 4 introduces words that we will use in section 5. We also prove Theorem 4.17 using this tool to illustrate how it works. Section 5 is devoted to the proof of Theorem 5.18. We start by explaining how to construct floor diagrams of genus $1$ from genus 00 floor diagrams. This allows us to express the genus $1$ asymptotic refined invariant as a weighted sum over the genus 00 floor diagrams. In both Sections 4 and 5 we start with the Hirzebruch case before going to the case of $h$-transverse non-singular toric surfaces.
@@ 2. Floor diagrams and asymptotic refined invariants
Let $N$ be a lattice and $M=\mathrm{Hom}(N,\mathbb{Z})$ its dual. We denote by $M_{\mathbb{R}}=M\otimes\mathbb{R}$, $N_{\mathbb{R}}=N\otimes\mathbb{R}$ the associated real vector spaces. Following [Ful93], a compact toric surface $X$ is obtained from a complete fan $\Sigma\subset N_{\mathbb{R}}$, or from a polygon $\Delta\in M_{\mathbb{R}}$ dual to $\Sigma$. The toric divisors of $X$ are in bijection with the rays of $\Sigma$. We assume that $X$ is smooth, i.e. every cone of $\Delta$ is simplicial. The anticanonical class $-K_{X}$ is represented by the sum of toric divisors. The Euler characteristic $\chi$ of $X$ is the equal to the number of rays of $\Sigma$.
@@
Consider the cone $\mathrm{Eff}(X)\subset H_{2}(X,\mathbb{Z})$ of effective classes, i.e. classes $\beta$ such that $\beta\cdot D>0$ for any toric divisor $D$. Let $\beta\in\mathrm{Eff}(X)$ be an effective class. For any ray $\rho$ of the fan $\Sigma$, let $n_{\rho}\in N$ be a primitive vector such that $\rho=\mathbb{R}_{\geqslant 0}\cdot n_{\rho}$, and let $D_{\rho}$ be the toric divisor corresponding to $\rho$. The multiset
@@
If $\beta\in\mathrm{Eff}(X)$, the lattice polygon $\Delta_{\beta}$ gives an ample line bundle $\mathcal{L}_{\beta}$ on $X$ with Chern class $c_{1}(\mathcal{L}_{\beta})\in H^{2}(X,\mathbb{Z})$ Poincaré dual to $\beta$. A basis of sections of $\mathcal{L}_{\beta}$ is indexed by the lattice points of $\Delta_{\beta}$. For our purpose, we work with the class $\beta$ instead of the line bundle, as it would be the case in the setting of the Göttsche conjecture. These points of view are equivalent in the case of rational surfaces. $\blacklozenge$
@@
so that $H_{2}(\mathbb{F}_{\delta},\mathbb{Z})$ is spanned by $F$ and $E$. The toric divisor associated to the last ray lies also in the class $F$. Let $\beta=aE+bF\in\mathrm{Eff}(\mathbb{F}_{\delta})$ be an effective class, which means that $a,b>0$. The associated tropical data is
@@
The corresponding polygon $\Delta_{aE+bF}$ is the trapezoid with vertices $(0,0)$, $(0,a)$, $(b,0)$ and $(b+\delta a,a)$, see Figure 1(b). $\lozenge$
@@ 2.2. Floor diagrams and tropical refined invariants
Let $\Delta$ be a $h$-transverse polygon. A floor diagram $\mathcal{D}$ with Newton polygon $\Delta$ is the data of $(\Gamma,w,\ell,r)$, with $(\Gamma,w)$ a weigthed, connected, oriented and acyclic graph satisfying the following conditions :
@@
Let $\mathcal{D}$ be a floor diagram of Newton polygon $\Delta$ and genus $g$. Its degree is the degree of the Laurent polynomial $(q^{1/2}-q^{-1/2})^{b^{\mathrm{top}}+b^{\mathrm{bot}}+2|E^{0}(\mathcal{D})|}\mu_{BG}(\mathcal{D})$:
@@
Let $\mathcal{D}$ be a genus $g$ marked diagram, and let $v_{1},\dots,v_{a}$ be the floors of $\mathcal{D}$, ordered by their marking but not necessarily inside $\mathcal{D}$. Let $\widetilde{\omega}_{m}$ be the sum of the weights of the bounded edges of $\mathcal{D}$ that link two floors $v_{m}$ and $v_{m+1}$. It may be 00 if there are no such edges. In comparison to a diagram of genus $g$ and codegree 0, the codegree of $\mathcal{D}$ comes from two different phenomena:
@@ 2.3. Asymptotic refined invariants
The first product is obtained clearing the denominators of each $[w]^{2}$. The second product comes from the ends: actually, each end contributes $[1]=\frac{q^{1/2}-q^{-1/2}}{q^{1/2}-q^{-1/2}}=1$ in $BG^{X}_{g}(\beta)$, and this becomes $q^{1/2}-q^{-1/2}$ when clearing denominators. The degree of the invariant $\overline{BG}^{X}_{g}(\beta)$ with cleared denominators is $\mathrm{Area}(\Delta_{\beta})$. This Laurent polynomial is symmetric (resp. antisymmetric) when $\beta\cdot K_{X}$ is even (resp. odd).
@@
$\begin{array}[]{crcl}\widetilde{BG}^{X}_{g}:&\mathrm{Eff}(X)&\longrightarrow&\mathbb{Z}[\![x]\!],\\ &\beta&\longmapsto&\widetilde{BG}^{X}_{g}(\beta)(x)\end{array}$
with values in the ring of formal series with integer coefficients $\mathbb{Z}[\![x]\!]$, even though for any $\beta$ the value is polynomial in $x$ of degree $2\mathrm{Area}(\Delta_{\beta})$. The codomain $\mathbb{Z}[\![x]\!]$ is a valuation ring, and can thus be endowed with the topology coming from the associated ultrametric distance. A basis of neighborhoods of 00 for this topology is given by the ideals $x^{n}\mathbb{Z}[\![x]\!]$, so that $f\in\mathbb{Z}[\![x]\!]$ is close to 00 if $f\equiv 0\ \mathrm{mod}\ x^{n}$ for $n\in\mathbb{N}$ sufficiently big. We prefer to use $\mathbb{Z}[\![x]\!]$ as codomain, because it is a complete space, more suited to express our asymptotic result. Meanwhile, we have a notion of neighborhood of infinity in the cone $\mathrm{Eff}(X)$: for $C>0$, we say that $\beta\succ C$ if $\beta\cdot D\geqslant C$ for every toric divisor $D$.
@@
$\widetilde{BG}^{X}_{g}(\beta)$ $=x^{\mathrm{Area}(\Delta_{\beta})}\displaystyle\sum_{(\mathcal{D},\mathfrak{m})}\ \displaystyle\prod_{e\in E^{0}(\mathcal{D})}(x^{-w(e)/2}-x^{w(e)/2})^{2}\displaystyle\prod_{e\in E^{\mathrm{top}}(\mathcal{D})\cup E^{\mathrm{bot}}(\mathcal{D})}(x^{-1/2}-x^{1/2})$
$=\displaystyle\sum_{(D,\mathfrak{m})}x^{\mathrm{codeg}(\mathcal{D})}\displaystyle\prod_{e\in E^{0}(\mathcal{D})}(1-x^{w(e)})^{2}\displaystyle\prod_{e\in E^{\mathrm{top}}(\mathcal{D})\cup E^{\mathrm{bot}}(\mathcal{D})}(1-x)$
$=\displaystyle\sum_{(D,\mathfrak{m})}\mu(\mathcal{D}).$
@@
The result amounts to prove that there exists polynomials $P^{X}_{g,i}(\beta)$ with degree bounded by a constant in $g$ such that we have
@@ 3.1. Degree 0
We start by computing $(AR^{X}_{g,0}$, the leading term of the asymptotic refined invariant. This amounts to compute the leading coefficient of the tropical refined invariant, which was already handled in [IM13, Proposition 2.11] using the lattice path algorithm from [Mik05]. We recall a proof here, because it uses a construction starting from $\mathcal{D}_{0}$ that will appear several times in subsection 3.2.
@@ 3.2. Degree 1
$\displaystyle\sum_{g\geqslant 0}AR_{g,1}^{X}u^{g}$ $=(1+u)^{g_{\max}}\left[-\beta^{2}\dfrac{u^{3}}{(1+u)^{3}}+2(K_{X}\cdot\beta)\dfrac{u^{2}}{(1+u)^{3}}+\chi\dfrac{1}{1+u}-K_{X}^{2}\dfrac{u}{(1+u)^{3}}\right].$
@@
The codegree 0 diagrams with Newton polygon $\Delta_{\beta}$ contribute to $\sum_{g}AR_{g,1}^{X}u^{g}$:
$(1+u)^{g_{\max}}\left[-(b^{\mathrm{top}}+b^{\mathrm{bot}})-2g_{\max}\frac{u^{2}}{(1+u)^{2}}-2(a-1)\frac{u(2+u)}{(1+u)^{2}}\right].$
@@
We construct a diagram of genus $g$ and codegree 0 as in the proof of Theorem 3.1. A diagram $\mathcal{D}$ of codegree 0 is counted with the degree 1 term of its multiplicity, that is
@@
The codegree 1 diagrams with an infinite side edge, with Newton polygon $\Delta_{\beta}$ contribute to $\sum_{g}AR_{g,1}^{X}u^{g}$:
$(1+u)^{g_{\max}}\left[(\omega_{1}+\omega_{a-1}-2)\frac{u}{(1+u)^{2}}+(b^{\mathrm{bot}}+b^{\mathrm{top}})\frac{1}{1+u}+2\frac{2+u}{(1+u)^{2}}\right].$
@@
The codegree 1 diagrams with a bounded side edge contribute to $\sum_{g}AR_{g,1}^{X}u^{g}$:
$(1+u)^{g_{\max}}\left[\displaystyle\sum_{j=1}^{a-2}\left(\omega_{j}+\omega_{j+1}-2\right)\frac{u^{2}}{(1+u)^{3}}+2(a-2)\frac{u(2+u)}{(1+u)^{3}}\right].$
@@
$\widetilde{\omega}_{j}=\omega_{j}-1,\ \widetilde{\omega}_{j+1}=\omega_{j+1}-1\text{ and }\widetilde{\omega}_{m}=\omega_{m},\ m\neq j.$
@@
The codegree 1 diagrams with an slope inversion contribute to $\sum_{g}AR_{g,1}^{X}u^{g}$:
$(\chi-4)(1+u)^{g_{\max}-1}.$
@@
$\displaystyle\sum_{g\geqslant 0}AR_{g,1}^{X}u^{g}$ $=(1+u)^{g_{\max}}\left[-\beta^{2}\dfrac{u^{3}}{(1+u)^{3}}+2(K_{X}\cdot\beta)\dfrac{u^{2}}{(1+u)^{3}}+\chi\dfrac{1}{1+u}-K_{X}^{2}\dfrac{u}{(1+u)^{3}}\right].$

### v2

@@ 1.1. Context
Given some points in the complex plane, the problem of determining how many curves of fixed degree and genus pass through these points is a well-known question, which generalizes to other surfaces. Given a non-singular complex surface $X$ equipped with a sufficiently ample line bundle $\mathcal{L}$, curves on $X$ may be obtained as zero-sets of sections of $\mathcal{L}$. For a non-negative integer $g$, how many such curves of genus $g$ pass through $c_{1}(X)\cdot\mathcal{L}-1+g$ points on $X$ We denote by $N_{g}^{X}(\mathcal{L})$ this number, which is also known as the degree of the corresponding Severi variety. As a consequence of the adjunction formula, we could consider a dual question: given $\delta$ a number of nodes, what is the number $N_{X}^{\delta}(\mathcal{L})$ of $\delta$-nodal curves passing through the appropriate number of points on $X$
@@ 1.1.3. Interpretations and applications of tropical refined invariants
    Bousseau proved [Bou19] that through the change of variable $q=e^{i\hbar}$, tropical refined invariants actually compute the generating series of (log-)Gromov Witten invariants with a $\lambda$-class.
@@ 1.2. Overview of results
For $X$ a smooth toric surface associated to a $h$-transverse and horizontal polygon, one has:
@@
We then give a result concerning generating series when summing over $i$: let $AR_{g}^{X}(\beta)=\sum_{i=0}^{\infty}AR^{X}_{g,i}(\beta)x^{i}$ be this series, which we call the genus $g$ asymptotic refined invariant.
In [Mév23], the second author proved that $AR^{X}_{0}(\beta)=p(x)^{\chi}$, where $p(x)$ is the generating series of the partition numbers. However, the method used there does not seem to be manageable to deal with higher genus. With a slightly different point of view, we give in Theorem 4.17 a different proof of this fact. The interest is that this point of view also allows to perform the computation in genus 1. A priori, the polynomiality behavior for general $h$-transverse toric surface has not been proven in [BJP22], but in the genus 0 and 1 case it actually follows from our computations.
@@
1.3. Future directions and conjectures
The form taken by the generating series for fixed genus suggests a nice but subtle regularity of the asymptotic of the refined invariants. However, we are for now limited by the computational techniques at our disposal, as the complexity of the computations increases quite fast with the genus or the codegree. We yet suggest a few possible generalizations that are either conjectures or potential results that the techniques hereby presented should be sufficient to prove for the brave reader but that we ultimately decided to leave out to avoid rendering this already quite technical paper even more technical.
First, although the results are stated for $h$-transverse polygons (especially the generating series for genus 00 and $1$), we expect them to be true for all polygons yielding smooth toric surfaces. To get that, one way is to relate the refined invariants of a toric surface $X$ and its blow-up at a torus fixed point, using floor diagrams or equivalently a version of the tropical Caporaso-Harris formula [GM07] near the corner/side corresponding to the torus fixed point/exceptional divisor. Although the Caporaso-Harris formula [GM07] is wrong in a general toric surface, we expect the latter to be true for sufficiently ample divisor classes in finite codegree.
[…]
With some time, the techniques presented should be able to provide a result for genus $2$ as well as codegree $2$, but at the cost of many lengthy computations. Given the above results, and that the result proven for $h$-transverse and non-singular toric surfaces should hold for any non-singular toric surface, we conjecture the following.
@@ 1.3. Future directions and conjectures
For non-singular surface $X$, the asymptotic refined invariant $AR^{X}_{g}$ has the following form:
@@
Conjecture 1.1 treats the regularity of the asymptotic invariant for a fixed surface $X$. Similarly to the Göttsche conjecture, it would be interesting to study the dependence of $AR_{g}^{X}$ in the surface $X$. Hopefully, the polynomials $AR_{g,i}^{X}(\beta)$ are actually given by a universal polynomial $Q_{g,i}(\beta^{2},\beta\cdot K_{X},K_{X}^{2},\chi(X))$. Theorems 3.2 and 5.18 prove that it is the case for $g=0,1$ or $i=0,1$.
@@ 1.4. Organization of the paper
Section 4 introduces words that we will use in section 5. We also prove Theorem 4.17 using this tool to illustrate how it works.
Section 5 is devoted to the proof of Theorem 5.18. We start by explaining how to construct floor diagrams of genus $1$ from genus 00 floor diagrams. This allows us to express the genus $1$ asymptotic refined invariant as a weighted sum over the genus 00 floor diagrams. In both Sections 4 and 5 we start with the Hirzebruch case before going to the case of $h$-transverse non-singular toric surfaces.
We end with an appendix where we explain how to modify the calculations of this paper to deal with Göttsche-Schroeter invariants instead of Block-Göttsche invariants.
@@ 2. Floor diagrams and asymptotic refined invariants
Let $N$ be a lattice and $M=\mathrm{Hom}(N,\mathbb{Z})$ its dual. We denote by $M_{\mathbb{R}}=M\otimes\mathbb{R}$, $N_{\mathbb{R}}=N\otimes\mathbb{R}$ the associated real vector spaces. Following [Ful93], a compact toric surface $X$ is obtained from a complete fan $\Sigma\subset N_{\mathbb{R}}$, or from a polygon $\Delta\in M_{\mathbb{R}}$ dual to $\Sigma$. The toric divisors of $X$ are in bijection with the rays of $\Sigma$. We assume that $X$ is smooth, i.e. every cone of $\Delta$ is simplicial. The anticanonical class $-K_{X}$ is represented by the sum of toric divisors. The Euler characteristic $\chi$ of $X$ is equal to the number of rays of $\Sigma$.
@@
Consider the cone $\mathrm{DAmp}(X)\subset H_{2}(X,\mathbb{Z})$ of ample divisors classes classes, i.e. classes $\beta$ such that $\beta\cdot D>0$ for any toric divisor $D$.
Let $\beta\in\mathrm{DAmp}(X)$ be an ample divisor class. For any ray $\rho$ of the fan $\Sigma$, let $n_{\rho}\in N$ be a primitive vector such that $\rho=\mathbb{R}_{\geqslant 0}\cdot n_{\rho}$, and let $D_{\rho}$ be the toric divisor corresponding to $\rho$. The multiset
@@
If $\beta\in\mathrm{DAmp}(X)$, the lattice polygon $\Delta_{\beta}$ gives an ample line bundle $\mathcal{L}_{\beta}$ on $X$ with Chern class $c_{1}(\mathcal{L}_{\beta})\in H^{2}(X,\mathbb{Z})$ Poincaré dual to $\beta$. A basis of sections of $\mathcal{L}_{\beta}$ is indexed by the lattice points of $\Delta_{\beta}$. For our purpose, we work with the class $\beta$ instead of the line bundle, as it would be the case in the setting of the Göttsche conjecture. These points of view are equivalent in the case of rational surfaces. $\blacklozenge$
@@
so that $H_{2}(\mathbb{F}_{\delta},\mathbb{Z})$ is spanned by $F$ and $E$. The toric divisor associated to the last ray lies also in the class $F$. Let $\beta=aE+bF\in\mathrm{DAmp}(\mathbb{F}_{\delta})$ be an ample divisor class, which means that $a,b>0$. The associated tropical data is
@@
The corresponding polygon $\Delta^{\delta}_{aE+bF}$ is the trapezoid with vertices $(0,0)$, $(0,a)$, $(b,0)$ and $(b+\delta a,a)$, see Figure 1(b). $\lozenge$
@@ 2.2. Floor diagrams and tropical refined invariants
Let $\Delta$ be a $h$-transverse polygon. A floor diagram $\mathcal{D}$ with Newton polygon $\Delta$ is the data of $(\Gamma,w,L,R)$, with $(\Gamma,w)$ a weigthed, connected, oriented and acyclic graph satisfying the following conditions :
@@
Let $\mathcal{D}$ be a floor diagram of Newton polygon $\Delta$ and genus $g$. We define its degree to be the degree of its multiplicity with cleared denominators, which is the Laurent polynomial $(q^{1/2}-q^{-1/2})^{b^{\mathrm{top}}+b^{\mathrm{bot}}+2|E^{0}(\mathcal{D})|}\mu_{BG}(\mathcal{D})$:
@@
Let $\mathcal{D}$ be a genus $g$ marked diagram, and let $v_{1},\dots,v_{a}$ be the floors of $\mathcal{D}$, ordered by their marking. Let $\widetilde{\omega}_{m}$ be the sum of the weights of the bounded edges of $\mathcal{D}$ that link two floors $v_{m}$ and $v_{m+1}$. It may be 00 if there are no such edges. In comparison to a diagram of genus $g$ and codegree 0, the codegree of $\mathcal{D}$ comes from two different phenomena:
@@ 2.3. Asymptotic refined invariants
The first product is obtained clearing the denominators of each $[w]^{2}$. The second product comes from the ends: actually, each end contributes $[1]=\frac{q^{1/2}-q^{-1/2}}{q^{1/2}-q^{-1/2}}=1$ in $BG^{X}_{g}(\beta)$, and this becomes $q^{1/2}-q^{-1/2}$ when clearing denominators.
The degree of the invariant $\overline{BG}^{X}_{g}(\beta)$ with cleared denominators is $\mathrm{Area}(\Delta_{\beta})$. This Laurent polynomial is symmetric (resp. antisymmetric) when $\beta\cdot K_{X}$ is even (resp. odd).
@@
$\begin{array}[]{crcl}\widetilde{BG}^{X}_{g}:&\mathrm{DAmp}(X)&\longrightarrow&\mathbb{Z}[\![x]\!],\\ &\beta&\longmapsto&\widetilde{BG}^{X}_{g}(\beta)(x)\end{array}$
with values in the ring of formal series with integer coefficients $\mathbb{Z}[\![x]\!]$, even though for any $\beta$ the value is polynomial in $x$ of degree $2\mathrm{Area}(\Delta_{\beta})$. The codomain $\mathbb{Z}[\![x]\!]$ is a valuation ring, and can thus be endowed with the topology coming from the associated ultrametric distance. A basis of neighborhoods of 00 for this topology is given by the ideals $x^{n}\mathbb{Z}[\![x]\!]$, so that $f\in\mathbb{Z}[\![x]\!]$ is close to 00 if $f\equiv 0\ \mathrm{mod}\ x^{n}$ for $n\in\mathbb{N}$ sufficiently big. We prefer to use $\mathbb{Z}[\![x]\!]$ as codomain, because it is a complete space, more suited to express our asymptotic result. Meanwhile, we have a notion of neighborhood of infinity in the cone $\mathrm{DAmp}(X)$: for $C>0$, we say that $\beta\succ C$ if $\beta\cdot D\geqslant C$ for every toric divisor $D$.
@@
$\widetilde{BG}^{X}_{g}(\beta)$ $=x^{\mathrm{Area}(\Delta_{\beta})}\displaystyle\sum_{(\mathcal{D},\mathfrak{m})}\ \displaystyle\prod_{e\in E^{0}(\mathcal{D})}(x^{-w(e)/2}-x^{w(e)/2})^{2}\displaystyle\prod_{\begin{subarray}{c}e\in E^{\mathrm{top}}(\mathcal{D})\\ \cup E^{\mathrm{bot}}(\mathcal{D})\end{subarray}}(x^{-1/2}-x^{1/2})$
$=\displaystyle\sum_{(\mathcal{D},\mathfrak{m})}x^{\mathrm{codeg}(\mathcal{D})}\displaystyle\prod_{e\in E^{0}(\mathcal{D})}(1-x^{w(e)})^{2}\displaystyle\prod_{\begin{subarray}{c}e\in E^{\mathrm{top}}(\mathcal{D})\\ \cup E^{\mathrm{bot}}(\mathcal{D})\end{subarray}}(1-x)$
$=\displaystyle\sum_{(\mathcal{D},\mathfrak{m})}\mu(\mathcal{D}).$
@@
The formulation of the result presented here amounts to prove that there exists polynomials $P^{X}_{g,i}(\beta)$ whose degree is bounded by a constant in $g$ such that we have
@@ 3.1. Degree 0
We start by computing $AR^{X}_{g,0}$, the leading term of the asymptotic refined invariant. This amounts to compute the leading coefficient of the tropical refined invariant, which was already handled in [IM13, Proposition 2.11] using the lattice path algorithm from [Mik05]. We recall a proof here, because it uses a construction starting from $\mathcal{D}_{0}$ that will appear several times in subsection 3.2.
@@ 3.2. Degree 1
The codegree 0 diagrams with Newton polygon $\Delta_{\beta}$ contribute
$(1+u)^{g_{\max}}\left[-(b^{\mathrm{top}}+b^{\mathrm{bot}})-2g_{\max}\frac{u^{2}}{(1+u)^{2}}-2(a-1)\frac{u(2+u)}{(1+u)^{2}}\right]$
to $\sum_{g}AR_{g,1}^{X}u^{g}$.
@@
We construct a diagram of genus $g$ and codegree 0 as in the proof of Proposition 3.1. A diagram $\mathcal{D}$ of codegree 0 is counted with the degree 1 term of its multiplicity, that is
@@
The codegree 1 diagrams with an infinite side edge, with Newton polygon $\Delta_{\beta}$ contribute
$(1+u)^{g_{\max}}\left[(\omega_{1}+\omega_{a-1}-2)\frac{u}{(1+u)^{2}}+(b^{\mathrm{bot}}+b^{\mathrm{top}})\frac{1}{1+u}+2\frac{2+u}{(1+u)^{2}}\right]$
to $\sum_{g}AR_{g,1}^{X}u^{g}$.
@@
The codegree 1 diagrams with a bounded side edge contribute
$(1+u)^{g_{\max}}\left[\displaystyle\sum_{j=1}^{a-2}\left(\omega_{j}+\omega_{j+1}-2\right)\frac{u^{2}}{(1+u)^{3}}+2(a-2)\frac{u(2+u)}{(1+u)^{3}}\right]$
to $\sum_{g}AR_{g,1}^{X}u^{g}$.
@@
$\widetilde{\omega}_{j}=\omega_{j}-1,\ \widetilde{\omega}_{j+1}=\omega_{j+1}-1\text{ and }\widetilde{\omega}_{m}=\omega_{m},\ m\notin\{j,j+1\}.$
@@
The codegree 1 diagrams with an slope inversion contribute
$(\chi-4)(1+u)^{g_{\max}-1}$
to $\sum_{g}AR_{g,1}^{X}u^{g}$.